# Leveraging Soft Prompts for Privacy Attacks in Federated Prompt Tuning

### v1

@@ 4.2 Performance of outlier detection methods against PromptMIA
Table 1: Outlier detection results averaged over 4 datasets and 3 models.
@@
Table 2, Figure 5 and Figure 12 show that naively applying anomaly detection is ineffective as a defense against PromptMIA. While IsolationForest achieves high recall, it does so by broadly flagging almost all benign keys as adversarial. LocalOutlierFactor and EllipticEnvelope detected almost none of the injected adversarial keys. OneClassSVM achieves moderate recall but also suffers from a high false positive rate. These findings highlight the ineffectiveness of traditional outlier detection methods against PromptMIA.
4.3 Performance and impact of noise perturbation against PromptMIA
@@
We evaluate the effectiveness of input noise perturbation as a defense against PromptMIA. In particular, we measure the attack success rate (ASR) of PromptMIA under three privacy budgets, $\epsilon\in\{3,5,8\}$, corresponding to high, medium, and low privacy levels, using the ViT-B32 model across four different datasets, as shown in Fig. 6. Table 2 reports the privacy–utility trade-off by presenting the model accuracy achieved under the same privacy settings. We find that using a larger $\delta_{\min}$ (0.20.20.30.3) works better under input noise perturbation. If $\delta_{\min}$ is too small, the adversarial keys are only slightly closer to $q(\mathcal{T})$ than the benign keys, so even a small amount of noise can break the attack. We also observe that input noise perturbation is largely ineffective at small batch sizes across all datasets: when the batch size is $1$ or $4$, the ASR remains close to $95\%$ even with the strongest privacy setting ($\epsilon=3$). As the batch size grows, increasing noise begins to reduce the attack’s success, but only when very strong privacy guarantees are applied. For instance, on CIFAR-100 with $\epsilon=3$, the ASR decreases noticeably at larger batch sizes, but this protection comes at a substantial cost in accuracy (dropping from $0.78$ to $0.50$). At moderate privacy levels ($\epsilon=5$ and $\epsilon=8$), input noise fails to offer meaningful protection: the attack continues to achieve moderate to high ASR on three out of four datasets, even when the batch size is large. These results highlight a privacy–utility trade-off: strong noise (small $\epsilon$) can partially suppress PromptMIA but severely harms accuracy, while weaker noise (larger $\epsilon$) preserves accuracy but leaves the model more vulnerable.
(a) CIFAR10
(b) CIFAR100
(c) TinyImagenet
(d) FourDataset
4.4 Ablation Experiments and Broader Discussion of Defense Strategies
@@
We analyze how key hyperparameters (global pool size $M$, selection size $N$, $\delta_{\min}$, $\Delta$, $\beta$, and number of training rounds) affect PromptMIA’s performance. The most interesting finding is that the attack success rate of PromptMIA is much higher on the models that have been trained for a few epochs rather than randomly initialized keys (Fig. 13(f)). This corresponds to our theoretical findings in Section 3.2.3 that FPR is lower when the benign data forms tight and compact clusters in the query space around the benign keys, which happens naturally during the training process ( Fig. 14). More detailed analysis and insights on hyperparameters are provided in Appendix A.9. We provide additional experiments on membership information leakeage under very large batch size in Appendix  A.11. To show that our attack generalizes to to all variants of federated prompt tuning that adopt the common paradigm of a frozen backbone model (often transformer-based) paired with a shared, learnable (soft) prompt pool across clients, we provide additional experiments on multimodal and text data in Appendix  A.12. Additional studies on the attack success rate and distribution of global keys and benign query vectors under heterogeneous settings is given in Appendix  A.13. Appendix A.14 presents a detailed discussion on the implications of other potential system-level defense strategies such as randomized key indices, secure aggregation Bonawitz et al. (2017), and prompt dropout. We show that these defenses are either provably or empirically ineffective.
@@ 5 Related Work
Membership Inference Attacks against Federated Learning. The goal of MIAs in FL is to identify if a specific data point was part of a client’s training set. The first AMI attack in FL was introduced by (Nasr et al., 2019), which required multiple FL iterations, while (Nguyen et al., 2023) proposed a stronger single-iteration AMI attack that relied on training a separate neural network. Both approaches incur non-trivial computational costs. (Vu et al., 2024) presented two low-complexity AMI attacks that exploit fully connected and attention layers in LLMs, achieving high success rates in compromising the membership privacy of unprotected client data. Zhu et al. (2025) proposed a three-step attack that leverage updates from all clients across multiple communication rounds.
Federated Fine-Tuning of Foundation Models. Federated fine-tuning avoids updating the entire model and instead allows clients to fine-tune only a small subset of parameters. These include LoRA-based methods (Qi et al., 2024; Wang et al., 2024; Fan et al., 2025), adapter-based approaches (Cai et al., 2022; Ghiasvand et al., 2024), selective-based tuning (Zaken et al., 2021; Yu et al., 2023). Recently, prompt-based federated fine-tuning approaches have been proposed (Su et al., 2024; Weng et al., 2024; Bai et al., 2024; Feng et al., 2024), which update soft prompts instead of full model weights during FL. While it is possible that LoRA and adapter-based federated fine-tuning may also exhibit similar vulnerabilities, these approaches are architecturally orthogonal to prompt-tuning.
Privacy Risks in LLM Prompting. LLMs are strong in-context learners that can adapt to downstream tasks by prepending discrete prompts such as exemplars or task instructions without requiring fine-tuning. These exemplars often contain sensitive information (e.g., medical records, personally identifiable data). Adversaries can exploit this by crafting malicious prompts to extract confidential information in these discrete prompts (Wen et al., 2024; Duan et al., 2023). To mitigate such risks, recent work has proposed a range of defense strategies primarily based on the notion of differential privacy (Duan et al., 2023; Wu et al., 2023; Hong et al., 2023; Tang et al., 2023).
@@ 6 Conclusion
In conclusion, our study reveals that federated prompt-tuning introduces a new and critical privacy vulnerability that cannot be easily addressed by existing defenses designed for gradient or output based attacks. Through PromptMIA, we demonstrate that a malicious server can exploit adversarial prompts to reliably infer client membership, achieving consistently high attack success rate across multiple datasets. Our theoretical analysis further explains the robustness of this attack by establishing a fundamental lower bound on its advantage. Finally, our evaluation of PromptMIA against MIA defenses underscores their limitations in this setting, highlighting the need for new defense strategies specifically tailored to the dynamics of federated prompt tuning.
Repoducibility Statement
We are committed to ensuring the reproducibility of our work. The code for this paper, along with trained model weights, will be released upon acceptance. All datasets and pretrained weights used in our experiments are publicly available. All hyperparameters are clearly specified in the paper. We also document software dependencies and hardware used for training and evaluation.
@@ A.2 Generating adversarial keys set $\mathcal{K}_{\textsc{adv}}$
1:Target query vector $q(\mathcal{T})\in\mathbb{R}^{D_{k}}$, desired cosine similarity $s\in(0,1)$
2:Adversarial key $k_{a}\in\mathbb{R}^{D_{k}}$ such that $\kappa(k_{a},q(\mathcal{T}))\approx s$ and $\|k_{a}\|=\|q(\mathcal{T})\|$
@@
1:Target query vector $q(\mathcal{T})\in\mathbb{R}^{D_{k}}$, benign key set $\mathcal{K}_{\textsc{benign}}$, margins $\delta_{\min},\Delta$, number of adversarial keys NN
@@
5:for $m=1$ to NN do
6:  Sample $s_{m}\sim\mathcal{U}(s_{\max}+\delta_{\min},\;s_{\max}+\delta_{\min}+\Delta)$
@@
8:end for
9:return Adversarial key set $\mathcal{K}_{\textsc{adv}}=\{k_{a_{m}}\}_{m=1}^{N}$
@@ A.3 Membership Inference Defenses
While the focus of our work is not on designing new defenses, we discuss how standard approaches—originally developed for gradient-based or output-based MIAs—can be adapted to, and interact with, PromptMIA.
@@
Noise Perturbation Rather than obfuscating gradient updates, clients can achieve differential privacy (DP) with respect to the input by injecting calibrated noise directly into the input image (Lecuyer et al., 2019). This introduces randomness to Eq. 1:
@@
Anomaly detection techniques We test commonly used anomaly detection techniques in machine learning against PromptMIA. Specifically, we consider the following classical approaches: IsolationForest (Liu et al., 2008), LocalOutlierFactor (Breunig et al., 2000), OneClassSVM (Manevitz and Yousef, 2001), SGDOneClassSVM, and EllipticEnvelope (Rosseeuw, 1999). These methods are widely used, well established in the literature, and directly available in scikit-learn library (Kramer, 2016). Although it would be interesting to evaluate deep learning–based anomaly detection algorithms (Do et al., 2025; Jiang et al., 2023) (e.g., autoencoders, VAEs, GAN-based models) against PromptMIA, we leave this for future work. Moreover, the number of prompts in the prompt pool is typically small (20 in our experiments), which raises doubts about the effectiveness of deep learning based methods in this setting. We therefore restrict our evaluation to the aforementioned classical methods. We briefly describe the anomaly detection algorithms use in this paper:
Isolation Forest detects anomalies by recursively partitioning the feature space with random splits. Each sample’s path length, or the number of splits needed to isolate it in a random tree—is shorter for outliers, as they are easier to separate from the bulk of the data. Averaging this path length over a forest of random trees yields an anomaly score: points with shorter average paths are more likely to be anomalous.
@@
A.4 Controlling the alignment of adversarial and benign keys
@@
1:Target query vector $q(\mathcal{T})\in\mathbb{R}^{D_{k}}$, desired cosine similarity $s\in(0,1)$, benign key set $\mathcal{K}_{\textsc{benign}}$, mixing factor $\beta\in(0,1)$
2:Vector $k_{a}\in\mathbb{R}^{D_{k}}$ such that $\kappa(k_{a},q(\mathcal{T}))\approx s$ and $\|k_{a}\|=\|q(\mathcal{T})\|$
@@
5:Sample $k_{b}\sim\mathcal{K}_{\textsc{benign}}$ $\triangleright$ Random benign key from the set
@@ A.5 Experimental Settings
We organize our evaluation along three main dimensions. First, we measure the performance of PromptMIA in terms of advantage ( Eq. 3) and attack success rate (Eq. 4) in Section 4.1. Second, we assess the robustness of PromptMIA against classical anomaly detection methods such as Isolation Forest, Local Outlier Factor, One-Class SVM, and Elliptic Envelope (Section 4.2). Third, we study the effect of noise-based defenses on the input image on PromptMIA attack performance (Section 4.3). Finally, we perform ablation experiments to analyze the impact of current number of training round, the role of the parameters $M$, NN, $\beta$, $\delta_{\min}$ and $\Delta$ on the performance of PromptMIA (Section 4.4).
@@
A.6 Proofs
Theorem 1 (True Positive Rate).
Let $\mathcal{K}_{\textsc{adv}}=\{k_{a_{m}}\}_{m=1}^{N}$ be the set of $N$ adversarial keys generated by Algorithm 2 with parameters $\delta_{\min}>0$ and $\Delta\geq 0$. Let $\mathcal{K}_{\textsc{benign}}$ be the set of MNM-N benign keys. If the client’s dataset $\mathcal{D}$ contains the target sample $\mathcal{T}$ (i.e., $b=1$) and the client’s selection mechanism (Eq. 1) selects the top-$N$ prompts based on highest cosine similarity (lowest cosine distance), the set of selected keys $\mathcal{K}_{\mathcal{T}}$ for the query $q(\mathcal{T})$ will be exactly the adversarial set:
$\mathcal{K}_{\mathcal{T}}=\mathcal{K}_{\textsc{adv}}$
Consequently, the True Positive Rate (TPR) of PromptMIA is 1.
@@ A.6 Proofs
Lemma 1 (Single-Point Flip Probability).
Let $q(x)\sim\mathcal{N}(k_{b,i},\sigma_{i}^{2}I)$ be a non-member query from benign cluster $i$. The probability $p_{i\to a}$ that $q(x)$ is closer to the adversarial key $k_{a}$ than to its own benign key $k_{b,i}$ is:
$p_{i\to a}=\operatorname{Pr}\left(\|q(x)-k_{a}\|^{2}<\|q(x)-k_{b,i}\|^{2}\right)=\Phi\left(-\frac{\|k_{b,i}-k_{a}\|}{2\sigma_{i}}\right)$
@@
The probability $p_{i\to a}$ is the probability of the ”flip” event, which is defined by:
$p_{i\to a}=\operatorname{Pr}\left(\|q(x)-k_{a}\|^{2}<\|q(x)-k_{b,i}\|^{2}\right)$
We first expand the squared Euclidean norms:
$\|q(x)-k_{a}\|^{2}<\|q(x)-k_{b,i}\|^{2}$
$\iff$ $(q(x)-k_{a})^{T}(q(x)-k_{a})<(q(x)-k_{b,i})^{T}(q(x)-k_{b,i})$
$\iff$ $q(x)^{T}q(x)-2k_{a}^{T}q(x)+\|k_{a}\|^{2}<q(x)^{T}q(x)-2k_{b,i}^{T}q(x)+\|k_{b,i}\|^{2}$
$\iff$ $-2k_{a}^{T}q(x)+\|k_{a}\|^{2}<-2k_{b,i}^{T}q(x)+\|k_{b,i}\|^{2}$
$\iff$ $2k_{b,i}^{T}q(x)-2k_{a}^{T}q(x)<\|k_{b,i}\|^{2}-\|k_{a}\|^{2}$
$\iff$ $2(k_{b,i}-k_{a})^{T}q(x)<\|k_{b,i}\|^{2}-\|k_{a}\|^{2}$
We set $Y=2(k_{b,i}-k_{a})^{T}q(x)$ and $C=\|k_{b,i}\|^{2}-\|k_{a}\|^{2}$. Based on Assumption 2, $q(x)$ is a multivariate Gaussian $q(x)\sim\mathcal{N}(k_{b,i},\sigma_{i}^{2}I)$. The variable $Y$ is a linear projection of $q(x)$, so $Y$ must also follow the normal Gaussian distribution (1D). We find its mean $E[Y]$ and variance $\text{Var}(Y)$:
@@
$E[Y]=E[(k_{b,i}-k_{a})^{T}q(x)]=(k_{b,i}-k_{a})^{T}E[q(x)]=(k_{b,i}-k_{a})^{T}k_{b,i}=\|k_{b,i}\|^{2}-k_{a}^{T}k_{b,i}$
Variance: We have $\text{Var}(Y)=\text{Var}((k_{b,i}-k_{a})^{T}q(x))$. Using $\text{Var}(Av)=A\text{Cov}(v)A^{T}$, with $A=(k_{b,i}-k_{a})^{T}$ and $\text{Cov}(q(x))=\sigma_{i}^{2}I$, we can infer that:
$\text{Var}(Y)=(k_{b,i}-k_{a})^{T}(\sigma_{i}^{2}I)(k_{b,i}-k_{a})$
$\text{Var}(Y)=\sigma_{i}^{2}(k_{b,i}-k_{a})^{T}(k_{b,i}-k_{a})=\sigma_{i}^{2}\|k_{b,i}-k_{a}\|^{2}$
Finally, we need to find $p_{i\to a}=\operatorname{Pr}(Y<C)$. We standardize $Y$ by converting it to $Z\sim\mathcal{N}(0,1)$:
$p_{i\to a}=\operatorname{Pr}\left(\frac{Y-E[Y]}{\sqrt{\text{Var}(Y)}}<\frac{C-E[Y]}{\sqrt{\text{Var}(Y)}}\right)=\Phi\left(\frac{C-E[Y]}{\sqrt{\text{Var}(Y)}}\right)$
The denominator is $\sqrt{\text{Var}(Y)}=\sigma_{i}\|k_{b,i}-k_{a}\|$. The numerator is CE[Y]C-E[Y]:
$C-E[Y]$ $=\left[\frac{1}{2}(\|k_{b,i}\|^{2}-\|k_{a}\|^{2})\right]-\left[\|k_{b,i}\|^{2}-k_{a}^{T}k_{b,i}\right]$
$=\frac{1}{2}\|k_{b,i}\|^{2}-\frac{1}{2}\|k_{a}\|^{2}-\|k_{b,i}\|^{2}+k_{a}^{T}k_{b,i}$
$=-\frac{1}{2}\|k_{b,i}\|^{2}+k_{a}^{T}k_{b,i}-\frac{1}{2}\|k_{a}\|^{2}$
$=-\frac{1}{2}\left(\|k_{b,i}\|^{2}-2k_{a}^{T}k_{b,i}+\|k_{a}\|^{2}\right)$
$=-\frac{1}{2}\left(k_{b,i}-k_{a}\right)^{T}\left(k_{b,i}-k_{a}\right)$
$=-\frac{1}{2}\|k_{b,i}-k_{a}\|^{2}$
We form the $z$-score by:
$z=\frac{-\frac{1}{2}\|k_{b,i}-k_{a}\|^{2}}{\sigma_{i}\|k_{b,i}-k_{a}\|}=-\frac{\|k_{b,i}-k_{a}\|}{2\sigma_{i}}$
Therefore, the probability is:
$p_{i\to a}=\Phi(z)=\Phi\left(-\frac{\|k_{b,i}-k_{a}\|}{2\sigma_{i}}\right)$
Theorem 2 (FPR Bound for $N=1$).
Let $\mathcal{B}$ be a client’s batch of $B$ non-member samples which includes $b_{i}$ samples from each benign cluster $i$ (i.e., $\sum_{i}b_{i}=B$). The False Positive Rate (FPR) of PromptMIA with $N=1$ is bounded by:
$FPR\leq\sum_{i=1}^{M-1}b_{i}\Phi\left(-\frac{D_{min}(i)}{2\sigma_{i}}\right)$
where $D_{min}(i)=\min_{c\in S}\|k_{b,i}-c\|$ is the minimum Euclidean distance from the benign key $k_{b,i}$ to the adversarial generation shell $S$.

### v2

@@ 4.2 Performance of Outlier Detection Dethods
Table 1: Outlier detection results averaged over all datasets and models.
@@
Table 1 shows that naively applying anomaly detection is ineffective against PromptMIA. While IsolationForest achieves high recall, it does so by broadly flagging almost all benign keys as adversarial. LocalOutlierFactor and EllipticEnvelope detected almost none of the injected adversarial keys. OneClassSVM achieves moderate recall but also suffers from a high false positive rate. These findings highlight the ineffectiveness of traditional outlier detection methods against PromptMIA. Figures 5 and 13 provide visualization of outlier detection methods against PromptMIA.
4.3 Performance and Impact of Noise Perturbation
@@
We evaluate the effectiveness of input noise perturbation as a defense against PromptMIA. In particular, we measure the attack success rate (ASR) of PromptMIA under three privacy budgets, $\epsilon\in\{3,5,8\}$, using the ViT-B32 model across four different datasets, as shown in Fig. 6. Table 2 reports the privacy-utility trade-off. We find that using a larger $\delta_{\min}$ ($0.2$) works better under input noise perturbation. If $\delta_{\min}$ is too small, the adversarial keys are only slightly closer to $q(\mathcal{T})$ than the benign keys, so even a small amount of noise can break the attack. As the batch size grows, increasing noise begins to reduce the attack’s success, but only when very strong privacy guarantees are applied. For instance, on CIFAR-100 with $\epsilon=3$, the ASR decreases noticeably at larger batch sizes, but this protection comes at a substantial cost in accuracy (dropping from $0.78$ to $0.50$). At moderate privacy levels ($\epsilon=5$ and $\epsilon=8$), input noise fails to offer meaningful protection: the attack continues to achieve moderate to high ASR on three out of four datasets, even when the batch size is large. These results highlight the trade-off: strong noise can partially suppress PromptMIA but severely harms accuracy, while weaker noise preserves accuracy but leaves the model more vulnerable.
4.4 Ablation Experiments
@@
We analyze how key hyperparameters (global pool size $M$, selection size $N$, $\delta_{\min}$, $\Delta$, $\beta$, and number of training rounds) affect PromptMIA’s performance. The most interesting finding is that the attack success rate of PromptMIA is much higher on the models that have been trained for a few epochs rather than randomly initialized keys (Fig. 14(f)). This corresponds to our theoretical findings in Section 3.2.3 that FPR is lower when the benign data forms tight and compact clusters in the query space around the benign keys, which happens naturally during the training process ( Appx. A.11). More detailed analysis and insights on hyperparameters are provided in Appendix A.10. We provide additional experiments under very large batch size in Appendix  A.12. To show that our attack generalizes to to all variants of FPT that adopt the common paradigm of a frozen backbone model (often transformer-based) paired with a shared, learnable (soft) prompt pool across clients, we provide additional experiments on multimodal and text data in Appendix  A.13. Additional studies on the attack success rate and distribution of global keys and benign query vectors under heterogeneous settings is given in Appendix  A.14.
@@ 5 Related Work
Membership Inference Attacks against Federated Learning. The goal of MIAs in FL is to identify if a specific data point was part of a client’s training set. Passive attacks (Shokri et al., 2017) involve an honest-but-curious server observing the model updates, while Active Membership Inference (AMI) attacks involve a dishonest server poisoning the global models, e.g., maliciously modifying model parameters, before dispatching them to clients (Nguyen et al., 2023; Vu et al., 2024). Recently, (Zhu et al., 2025) proposed a three-step attack that leverages updates from all clients that can be integrated as an extension to existing attacks.
Federated Fine-Tuning of Foundation Models. Federated fine-tuning avoids updating the entire model and instead allows clients to fine-tune only a small subset of parameters. These include LoRA-based methods (Qi et al., 2024; Wang et al., 2024; Fan et al., 2025), adapter-based approaches (Cai et al., 2022; Ghiasvand et al., 2024). Recently, prompt-based federated fine-tuning approaches have been proposed (Su et al., 2024; Weng et al., 2024; Bai et al., 2024; Feng et al., 2024), which update soft prompts instead of full model weights during FL. While it is possible that LoRA and adapter-based federated fine-tuning may also exhibit similar vulnerabilities, our work focuses on MIAs under FPT setting since these approaches are architecturally orthogonal.
Privacy Risks in LLM Prompting. LLMs are strong in-context learners that can adapt to downstream tasks by prepending discrete prompts such as exemplars or task instructions without requiring fine-tuning. These exemplars often contain sensitive information (e.g., medical records, personally identifiable data). Adversaries can exploit this by crafting malicious prompts to extract confidential information in these discrete prompts (Wen et al., 2024; Duan et al., 2023). To mitigate such risks, recent work has proposed a range of defense strategies primarily based on the notion of differential privacy (Duan et al., 2023; Wu et al., 2023; Hong et al., 2023; Tang et al., 2023). Our work is the first to investigate the privacy risks of soft prompts in FPT.
@@ 6 Conclusion
In conclusion, we show that FPT introduces a new and critical privacy vulnerability. Through PromptMIA, we demonstrate that a malicious server can leverage adversarial prompts to reliably infer client membership, achieving high attack success rates across multiple datasets. Our theoretical analysis explains the attack’s robustness by establishing a lower bound on the attack’s advantage. Finally, evaluation of PromptMIA against existing MIA defenses reveals their limitations in this setting, underscoring the need for new defense strategies specifically designed for FPT.
Impact Statement
This work reveals that federated prompt tuning introduces new privacy risks, showing that active membership inference attacks can compromise client data under practical federated learning settings. Our results demonstrate that existing defenses are either ineffective or incur prohibitive utility costs to provide meaningful protection. These findings have important implications for deploying federated prompt-tuning in privacy-sensitive domains such as healthcare and finance, underscoring the need for prompt-aware privacy safeguards and more robust defenses. We hope this work motivates further research on secure and privacy-preserving prompt-based federated learning.
@@ A.2 Generating adversarial keys set $\mathcal{K}_{\textsc{adv}}$
0: Target query vector $q(\mathcal{T})\in\mathbb{R}^{D_{k}}$, desired cosine similarity $s\in(0,1)$
0: Adversarial key $k_{a}\in\mathbb{R}^{D_{k}}$ such that $\kappa(k_{a},q(\mathcal{T}))\approx s$ and $\|k_{a}\|=\|q(\mathcal{T})\|$
@@
0: Target query vector $q(\mathcal{T})\in\mathbb{R}^{D_{k}}$, benign key set $\mathcal{K}_{\textsc{benign}}$, margins $\delta_{\min},\Delta$, number of adversarial keys NN
@@
3:for $m=1$ to NN do
4:  Sample $s_{m}\sim\mathcal{U}(s_{\max}+\delta_{\min},\;s_{\max}+\delta_{\min}+\Delta)$
@@
6:end for
7:return Adversarial key set $\mathcal{K}_{\textsc{adv}}=\{k_{a_{m}}\}_{m=1}^{N}$
@@ A.3 Membership Inference Defenses
While the focus of our work is not on designing new defenses, we discuss how standard approaches originally developed for gradient-based or output-based MIAs can be adapted to, and interact with, PromptMIA.
@@
Noise Perturbation Rather than obfuscating gradient updates, clients can achieve differential privacy (DP) w.r.t the input by injecting calibrated noise directly into the input image (Lecuyer et al., 2019). This introduces randomness to Eq. 1:
@@
Anomaly detection techniques We test commonly used anomaly detection techniques in machine learning against PromptMIA. Specifically, we consider the following classical approaches: IsolationForest (Liu et al., 2008), LocalOutlierFactor (Breunig et al., 2000), OneClassSVM (Manevitz and Yousef, 2001), and EllipticEnvelope (Rosseeuw, 1999). These methods are widely used, well established in the literature, and directly available in scikit-learn library (Kramer, 2016). Although it would be interesting to evaluate deep learning-based anomaly detection algorithms (Do et al., 2025; Jiang et al., 2023) (e.g., autoencoders, VAEs, GAN-based models) against PromptMIA, we leave this for future work. Moreover, the number of prompts in the prompt pool is typically small (20 in our experiments), which raises doubts about the effectiveness of deep learning based methods in this setting. We therefore restrict our evaluation to the aforementioned classical methods. We briefly describe the anomaly detection algorithms used in this paper:
  • Isolation Forest detects anomalies by recursively partitioning the feature space with random splits. Each sample’s path length, or the number of splits needed to isolate it in a random tree is shorter for outliers, as they are easier to separate from the bulk of the data. Averaging this path length over a forest of random trees yields an anomaly score: points with shorter average paths are more likely to be anomalous.
@@
Figure 8: Impact of Prompt Dropout on attack success rate.
A.4 System-level Defenses
We also analyze the impact of system-level defenses that modify the prompt selection or aggregation mechanism against PromptMIA.
  • Secure Aggregation. Another strategy is to apply secure aggregation (Bonawitz et al., 2017) so that individual client updates are hidden from the server. The intuition is that, if the server only observes an aggregated result rather than each client’s prompt updates, it may be unable to determine whether a particular adversarial prompt was selected. However, current secure aggregation protocols are developed for linear aggregation (that is, weighted averaging) of local model parameters, and do not extend to probablistic aggregation methods commonly used in federated prompt tuning. Weighted averaging is not sufficient in non-IID settings where local parameters must first be aligned before aggregation in order to prevent important features from collapsing into less informative representations due to semantic misalignment (Weng et al., 2024). To address this, the recent PFPT work (Weng et al., 2024) developed an aligned and aggregated mechanism based on probabilistic non-parametric clustering which was shown to achieve substantially better performance than weighted averaging in non-IID settings. However, both the alignment and the aggregation steps in PFPT are inherently non-linear, making it unclear how existing secure aggregation methods could be generalized to support such a mechanism.
  • Prompt Dropout. Since PromptMIA relies on all adversarial keys being selected, introducing randomness into the prompt selection (e.g via prompt dropout) can reduce the attack success rate but it remains substantial (larger than 75%) as shown in our experiment with prompt dropout in Fig. 8.
A.5 Controlling the alignment of adversarial and benign keys
@@
0: Target query vector $q(\mathcal{T})\in\mathbb{R}^{D_{k}}$, desired cosine similarity $s\in(0,1)$, benign key set $\mathcal{K}_{\textsc{benign}}$, mixing factor $\beta\in(0,1)$
0: Vector $k_{a}\in\mathbb{R}^{D_{k}}$ such that $\kappa(k_{a},q(\mathcal{T}))\approx s$ and $\|k_{a}\|=\|q(\mathcal{T})\|$
@@
3:Sample $k_{b}\sim\mathcal{K}_{\textsc{benign}}$ {Random benign key from the set}
@@ A.6 Experimental Settings
We organize our evaluation along three main dimensions. First, we measure the performance of PromptMIA in terms of advantage ( Eq. 3) and attack success rate (Eq. 4) in Section 4.1. Second, we assess the robustness of PromptMIA against classical anomaly detection methods such as Isolation Forest, Local Outlier Factor, One-Class SVM, and Elliptic Envelope (Section 4.2). Third, we study the effect of noise-based defenses on the input image on PromptMIA attack performance (Section 4.3). Finally, we perform ablation experiments to analyze the impact of current number of training round, data heterogeineity, input modalities and the role of the parameters $M$, NN, $\beta$, $\delta_{\min}$ and $\Delta$ on the performance of PromptMIA (Section 4.4).
@@
A.7 Proofs
Theorem 3.1 (True Positive Rate).
@@ A.7 Proofs
Lemma 3.2.
@@
Let $E_{i}$ be the event that a query $q(x)$, drawn from the $i$-th benign cluster (i.e., $q(x)\sim\mathcal{N}(k_{b_{i}},\sigma_{i}^{2}I)$), selects all $N$ adversarial keys. This occurs when all adversarial keys have a higher similarity to $q(x)$ than all MNM-N benign keys. This is formally defined as:
$E_{i}=\left\{\min_{j=1...N}q(x)^{T}k_{a_{j}}>\max_{l=1...M-N}q(x)^{T}k_{b_{l}}\right\}$
Here, we omit the normalization factors, since both the query and all keys are $\ell_{2}$-normalized. This is an intersection of $N\times(M-N)$ pairwise events, $E_{i}=\bigcap_{j=1}^{N}\bigcap_{l=1}^{M-N}A_{jl}$, where:
$A_{jl}=\left\{q(x)^{T}k_{a_{j}}>q(x)^{T}k_{b_{l}}\right\}\iff A_{jl}=\left\{(k_{a_{j}}-k_{b_{l}})^{T}q(x)>0\right\}$
The probability of an intersection of events is less than or equal to the probability of the single least likely event in that set. This gives us a strict upper bound:
$\Pr(E_{i})=\Pr\left(\bigcap_{j,l}A_{jl}\right)\leq\min_{\begin{subarray}{c}1\leq j\leq N\\ 1\leq l\leq M-N\end{subarray}}\Pr(A_{jl})$
Now, we find the probability of a single pairwise event $A_{jl}$. Let $Y_{jl}=(k_{a_{j}}-k_{b_{l}})^{T}q(x)$. Based on Assumption 2, $q(x)$ is a multivariate Gaussian $q(x)\sim\mathcal{N}(k_{b_{i}},\sigma_{i}^{2}I)$. The variable $Y_{jl}$ is a linear projection of $q(x)$, so it also follows a 1D Gaussian distribution. We find its mean $E[Y_{jl}]$ and variance $\text{Var}[Y_{jl}]$:
@@
$E[Y_{jl}]$ $=E[(k_{a_{j}}-k_{b_{l}})^{T}q(x)]=(k_{a_{j}}-k_{b_{l}})^{T}E[q(x)]=(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}$
Variance:
$\text{Var}[Y_{jl}]$ $=\text{Var}((k_{a_{j}}-k_{b_{l}})^{T}q(x))=(k_{a_{j}}-k_{b_{l}})^{T}\text{Cov}(q(x))(k_{a_{j}}-k_{b_{l}})$
$=(k_{a_{j}}-k_{b_{l}})^{T}(\sigma_{i}^{2}I)(k_{a_{j}}-k_{b_{l}})=\sigma_{i}^{2}\|k_{a_{j}}-k_{b_{l}}\|^{2}$
We want to find $\Pr(A_{jl})=\Pr(Y_{jl}>0)$. We standardize $Y_{jl}$ to $Z\sim\mathcal{N}(0,1)$ as:
$\Pr(Y_{jl}>0)=\Pr\left(Z>\frac{0-E[Y_{jl}]}{\sqrt{\text{Var}(Y_{jl})}}\right)=\Pr\left(Z>-\frac{E[Y_{jl}]}{\sqrt{\text{Var}(Y_{jl})}}\right)$
Using the identity $\Pr(Z>-z)=\Phi(z)$, the probability is $\Phi(z_{ijl}(\mathcal{K}_{ADV}))$, where the $z$-score is:
$z_{ijl}(\mathcal{K}_{ADV})=\frac{E[Y_{jl}]}{\sqrt{\text{Var}(Y_{jl})}}=\frac{(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}}{\sigma_{i}\|k_{a_{j}}-k_{b_{l}}\|}$
Substituting this result into the original inequality, we have the final bound:
Theorem 3.3 (False Positive Rate).
The per-sample False Positive Rate (FPR) is bounded by:
where $\mathcal{K}_{ADV}=\{k_{a_{1}},\dots,k_{a_{N}}\}$ is the set of $N$ adversarial keys drawn from the shell $S$, and $z_{ijl}$ is the $z$-score:
$z_{ijl}=\frac{(k_{a_{j}}-k_{b_{l}})^{T}k_{b_{i}}}{\sigma_{i}\|k_{a_{j}}-k_{b_{l}}\|}$